# TransVIP: Speech to Speech Translation System with Voice and Isochrony Preservation

### v1

@@ Abstract
There is a rising interest and trend in research towards directly translating speech from one language to another, known as end-to-end speech-to-speech translation. However, most end-to-end models struggle to outperform cascade models, i.e., a pipeline framework by concatenating speech recognition, machine translation and text-to-speech models. The primary challenges stem from the inherent complexities involved in direct translation tasks and the scarcity of data. In this study, we introduce a novel model framework TransVIP that leverages diverse datasets in a cascade fashion yet facilitates end-to-end inference through joint probability. Furthermore, we propose two separated encoders to preserve the speaker’s voice characteristics and isochrony from the source speech during the translation process, making it highly suitable for scenarios such as video dubbing. Our experiments on the French-English language pair demonstrate that our model outperforms the current state-of-the-art speech-to-speech translation model.
Figure 1: Overview of our speech to speech translation framework: 1) Joint encoder-decoder model for translating speech into target text, and coarse-grained speech tokens, $C_{0}$; 2) Non-autoregressive acoustic model for acoustic details, $C_{0:16}$; 3) Codec model to convert discrete speech tokens back to waveform. Abbreviation: S/A/I(Semantic/Acoustic/Isochrony Information), $C_{0}$/$C_{0:16}$(Codec layer 0/0-15), S/A-Enc(Semantic/Acoustic Encoder), ICM(Isochrony Control Module).
@@ 1 Introduction
To address these issues, we propose TransVIP, a speech-to-speech Translation framework with Voice and Isochrony Preservation. TransVIP employs a consecutive generation approach, simplifying the complex S2ST task into two sequential tasks while maintaining an end-to-end framework. The generation of the target speech is conditioned not only on the semantic information, as in conventional S2ST models, but also on the isochrony and acoustic information derived from the source speech. The corresponding overview of TransVIP is demonstrated in Figure 1. We evaluate the performance of TransVIP for French-English mutual translation using a subset of the CVSS-T test set [8]. The results demonstrate that TransVIP outperforms the publicly available SOTA models such as a larger Seamless Expressive model. The generated audio and video samples are available in https://aka.ms/transvip . The main contributions of the paper can be summarized as follows:
@@
    We introduce a framework for speech-to-speech translation tasks that employs a consecutive generation with joint inference. This method efficiently utilizes a variety of datasets through multi-task learning to overcome the challenge of scarce paired data during the training phase, while preserving the end-to-end nature during inference.
@@
    We propose to disentangle various information required to learn in the training stage through employing separated encoders. It can enhance the transfer of voice characteristics and isochrony/temporal-alignment from the source to the target speech in the translation process. Additionally, it facilitates the design of lightweight modules for more effective individual information learning.
@@ 2 Related Works
The task of speech quantization is to transform continuous speech feature into discrete tokens. The quantization module is usually trained by self-supervised learning (SSL) methods. The speech tokens can be divided into two categories: semantic tokens and acoustic tokens[11]. The semantic tokens[12, 13, 14, 15, 16, 17] are usually learned by context prediction task. This kind of tokens is rich in context information and is suitable for downstream task like recognition. The acoustic tokens[18, 19, 20, 21, 22], also called the neural codec, is usually trained by reconstruction task through a discrete information bottleneck. This codec is suitable for audio compression and audio generation. SpeechTokenizer [9] combines the semantic token and acoustic token by semantic distillation. Latest works propose attribute disentanglement of neural codec for better generation task support. FAcodec[23] further disentangle codec into content, prosody, acoustic and timbre through supervision and reversed gradient.
@@
With the recent advance in speech quantization and the strong in-context learning capability of language model, speech language model has shown great potential in speech generation. The GSLM family[24, 25, 26] use semantic tokens to train language model on speech continuation. VoxtLM [27] integrates text tokens with semantic tokens to perform speech recognition, speech synthesis, text generation, and speech continuation in a single causal language model. AudioLM[11] use both semantic tokens and acoustic tokens for high-quality speech continuation. VALL-E [10] extends this idea to zero-shot TTS and proposes auto-regressively generating the first layer and non-autoregressively the rest layers of the codec codes. Later work[28, 29, 30] have enhanced the functionality of zero-shot TTS systems, such as building cross-lingual synthesis and accurate alignment between text control signal and acoustic tokens.
@@
An End-to-end approach has been an ultimate goal for S2ST research. Some of the previous works[31, 32] have tried direct end-to-end approach for S2ST. However, by far such methods still suffer great performance issue. A solution is to include semantic tokens(text, phoneme, or semantic speech tokens) as a intermediate result. Some recent works[33, 6, 34, 35] applied speech-to-unit translation(S2UT) that translate speech input into semantic speech tokens and then generate speech through a unit vocoder like hifi-gan. And UnitY[36] use separate speech to text and text to unit models that can be optimized together. However traditional vocoder based on semantic tokens like hifi-gan cannot generate high-quality speech. SeamlessExpressive[37] address this issue using a well-designed PRETSSEL vocoder. PolyVoice[38] and AudioPaLM[39] employees use two causal model to first translate speech into semantic tokens, then generate acoustic tokens based on semantic tokens and then convert acoustic tokens to output speech. VioLA[40] and MSLM-S2ST[41] adopts a similar approach but use a single causal model in a multi-tasking way. However, these approaches perform several separate auto-regressive generations during the inference process. This, to a certain degree, contradicts the principle of end-to-end processing.
Figure 2: The illustration of the training framework of the Joint Enc-Dec Model. During the training, the losses from the target speech clip, i.e., a sub-part of the whole target speech, which serves as a prompt, are not aggregated when computing the Cross-Entropy (CE) loss. The corresponding codec labels are masked in the implementation. The semantic encoder and the auto-regressive decoder are initialized by a SeamlessM4T X2T model  [^2] https://ai.meta.com/blog/seamless-m4t. The semantic encoder are frozen during training. In inference, all the target speech input are replaced by source speech input.
@@ 3 TransVIP
In this section, we present the design details of our TransVIP system. Inspired by the recent advance in codec-based zero-shot TTS such as VALL-E [10], TransVIP contains three major part: 1) A codec model for speech quantization and reconstruction, which is essential for auto-regressive speech generation (Section 3.2). 2) An auto-regressive joint translation model to translate the input speech in to coarse-grained speech (Section 3.1). 3) A textless non-autoregressive acoustic model to replenish acoustic details into the output speech (Section 3.3). The overall framework for generating target speech is shown in Figure 1.
3.1 Joint Translation with Multiple Controls
In this section we will present a detailed introduction to our joint translation model, which is the key component of our translation system. We call it a joint model owning to its support for multiple input modal (speech/text), multiple output modal (speech/text), multiple control signal (voice/isochrony) and a joint probability modeling for text and speech generation. The training procedure of the joint model is illustrated in Figure 2.
3.1.1 Consecutive Generation with Joint Inference
Directly generating target translation speech $Y$ from input speech $X$ presents a considerable challenge. Therefore, the primary objective of this model is to produce a coarse-grained speech $Y^{\prime}$ represented by a first-layer codec sequence. Instead of optimizing the direct probability distribution of $P(Y^{\prime}|X)$, our optimizing target is the joint probability $P(Y^{\prime},T|X)$ , in order to preserve semantic accuracy in the text outputs alone, as $P(T|X)$. Hence, our approach focuses on modeling and maximizing the conditional expression:
@@ 3.1.1 Consecutive Generation with Joint Inference
3.1.2 Feature Disentanglement
In the real world, an ideal training dataset for speech-to-speech generation, where the source and target speeches are roughly equal in length and uttered by the same speaker, is not readily available. A real dataset is composed of paired speech which is either spoken by different speakers or synthesized. Thus, we cannot guarantee voice and duration preservation with these corpus.
To address this issue, we decouple the input feature into three parts: semantic information($S$), acoustic information($A$) and isochrony information($I$). Formally, $X=(S,A,I)$. During training the inputs to the model are $S$ from source speech, and $(A,I)$ information from reference target speech. While during inference, $(A,I)$ are derived from source speech. In this way the input feature can always match the output speech in terms of voice characterises $A$ and isochrony $I$.
@@ 3.1.2 Feature Disentanglement
$S$ is extracted using a pretrained speech encoder from an encoder-decoder speech/text-to-text translation model, which is designed to provide minimal acoustic or speaker information. We also include the text encoder of the pretrained model to enable text input and perform a knowledge distillation loss on the text output part. We freeze two pretrained encoders during the training process.
@@
The acoustic information that we extract relate to the distinctive acoustic characteristics of a speaker’s voice. We assume that the translated text $T$ is independent to $A$. The relationship can be modeled as in Equation 3. And we take advantage of the one-direction mask in auto-regressive decoder to explicitly model such relation.
@@
We use an acoustic encoder, comprising transformer blocks that learn from scratch, to extract the acoustic information. This information is sum pooled along the time dimension into a single embedding, which then replaces the embedding of a separation token at the input of the auto-regressive decoder. This setup prevents text tokens from attending to the acoustic embedding, and is compatible with both gradient passing during training and beam search during inference. Similar to the idea of Classifier-free Guidance, this replacement operation is not performed with a certain probability(in our experiment $p=0.5$) during training. Moreover, during training, the input to the acoustic encoder is a speech prompt, i.e., a sub-part of the whole target speech. To prevent the model from directly passing the speech information, we zero out the loss in the corresponding part of the label.
@@
We define a frame length, in our case 160ms, to quantify the duration of speech. Our Isochrony control module contains three embeddings: a regular position embedding, a reversed position embedding, and a voice activity detection (VAD) 0-1 embedding. The reversed position embedding provides the model with duration information on the number of additional tokens to be generated in each step of the inference process [42], and the VAD embedding identifies whether each frame is voice-active or not. Isochrony information is then represented by sum of the three embedding sequences, which is then concatenated with the semantic feature $S$ before fed into the decoder through cross attention. This setup enables our model to align with both the total length and the voice-active regions of the input speech, which is especially beneficial for translating long speeches with pauses.
3.1.3 Multi-task Training
@@
  • S2ST Dataset comprises quadruple of ($X$, $T_{s}$, $T_{t}$, $Y$), where $T_{s}$ is the source text and $T_{t}$ is the target text. We either use $X$ or $T_{s}$ as input and consecutive $T_{t}$ and $Y$ as target label, or we can reverse the roles of source input and target label.
  • ST Dataset is a triplet of ($X$, $T_{s}$, $T_{t}$). In our model the missing of codec label does not impede the training of speech-to-text translation. We can use $X$ as the input and treat $T_{t}$ as an incomplete label that terminates with a separation token. In this setup, the isochrony feature is omitted and does not concatenate with the semantic feature. And, alternatively, $T_{t}$ can serve as the input with consecutive $X$ and $T_{s}$ as the label, mirroring the approach in the S2ST dataset.
  • ASR Dataset, which is highly accessible, consists of pairs ($X$, $T_{s}$). Unlike many other speech translation work, we use ASR data mainly for TTS training, fitting $P(Y^{\prime}|S,A,I,T)$. Here, $T_{s}$ is used as the input with consecutive $T_{s}$ and $X$ as the output labels, and the loss is not applied to the part of the label corresponding to $T_{s}$. We didn’t train on recognition task because our semantic encoder is frozen, making this task ineffective.
@@ 3.2 Nerual Codec with Semantic Distillation
The neural codec we utilize is designed to provide not only high reconstruction quality but also compatibility with language model predictions. We refer to our codec as the multi-lingual Semantic-Aware Speech Codec (SASC), which integrates two existing methodologies: DAC [22] and SpeechTokenizer. DAC employs innovative factorized and L2-normalized codes in vector quantization, along with finely tuned hyperparameters, to achieve state-of-the-art reconstruction quality. SpeechTokenizer, on the other hand, uses a Hubert encoder as a teacher and performs semantic distillation between it and the first-layer Residual Vector Quantization (RVQ) codes. This process enriches the RVQ codes with additional semantic information, thereby facilitating downstream language modeling tasks.
@@
To enhance reconstruction quality, we have integrated the codec with factorized and L2-normalized codes. The discriminator, loss configuration, and training hyperparameters are all aligned with those used in DAC.
@@ 3.3 Textless Non-autoregressive Acoustic Modeling
Similar to previous work, the function of our non-autoregressive model is to enhance the acoustic details of generated speech by predicting the subsequent layers of Residual Vector Quantization (RVQ) codes based on the initial layer. We utilize a standard transformer equipped with adaptive layer normalization for this task. The model inputs the summation of embeddings from the first $n$ layers of RVQ codes to predict the $(n+1)$th layer, where $n=1,2,...,N$ and $N$ represents the total number of RVQ layers. A prompt is concatenated with the input embeddings along time dimension. The prompt is the summation of full $N$ layers of embeddings of the prompt speech. During the training phase, both the prompt and the part to be generated are in the same language. However, in the inference phase, they are in two distinct languages for the translation task. This creates a certain gap. Fortunately, our model trained using a multilingual dataset, is capable of generalizing and effectively bridging this gap.
Our approach diverges from prior studies by excluding any textual or phonetic inputs, rendering it a completely textless model. Previous research indicated that omitting semantic information such as text or phoneme typically results in a significant decrease in intelligibility. However, our experiments demonstrate that the textless model either matches or surpasses the performance of models with text inputs, owing to semantic distillation within the SASC. An comparison of text and textless model is shown in appendix B.3
The textless model offers two significant advantages: Firstly, it improves data accessibility, as the acoustic model can now be trained entirely with unsupervised data. Unlike supervised data, which requires aligned text labels and precise speech clip boundaries, unsupervised data can be arbitrarily segmented, effectively augmenting the dataset. Secondly, traditional non-autoregressive models that utilize phoneme inputs depend on accurate transcriptions to establish alignment and initiate generation. This requirement can be challenging, especially in zero-shot speech translation tasks where users might not provide transcriptions. Our textless model addresses this problem by completely eliminating the need for transcription.
@@
In previous work, the decoding of non-autoregressive models is typically performed greedily, as the decoding techniques used in auto-regressive models cannot be directly applied. Greedy decoding often leads to a problem known as "early decision error," where the model is unable to revise previous outputs once generated. To address this, we propose a method called Layer Beam Search (LBS), which adapts the beam search algorithm for use in non-autoregressive models.
Consider a beam size of $B$ and a vocabulary size of $V$. In a typical beam search algorithm, a beam maintains $B$ hypotheses. At each decoding step, each hypothesis calculates the probability distribution over $V$, yielding $B\times V$ candidates. These candidates are then ranked by their aggregated scores, and the top $B$ are selected to form the new beam. This approach is impractical in non-autoregressive models due to the excessively large vocabulary size at each decoding step. For example, for a speech segment containing $L$ tokens, and a codec codebook size of $C$, the effective vocabulary size becomes $|V|=C^{L}$, which is too large to enumerate.
To tackle this, we employ a sampling method to choose a reasonable number of candidates. Rather than sampling from the entire vocabulary space $V$, we sample from the top-K candidates at each token, thus creating a reduced vocabulary subspace $V^{\prime}$ of size $k^{L}$. We sample $N$ times from $V^{\prime}$ for each candidate in original beam to form in total $B*N$ new candidates. We then rank and form a new beam of size $B$ from these sampled candidates. The pseudo code for the Layer Beam Search is provided in Appendix A.

### v2

@@ Abstract
There is a rising interest and trend in research towards directly translating speech from one language to another, known as end-to-end speech-to-speech translation. However, most end-to-end models struggle to outperform cascade models, i.e., a pipeline framework by concatenating speech recognition, machine translation, and text-to-speech models. The primary challenges stem from the inherent complexities involved in direct translation tasks and the scarcity of data. In this study, we introduce a novel model framework TransVIP that leverages diverse datasets in a cascade fashion yet facilitates end-to-end inference through joint probability. Furthermore, we propose two separate encoders to preserve the speaker’s voice characteristics and isochrony from the source speech during the translation process, making it highly suitable for scenarios such as video dubbing. Our experiments on the French-English language pair demonstrate that our model outperforms the current state-of-the-art speech-to-speech translation model.
Figure 1: Overview of our speech-to-speech translation framework: 1) Joint encoder-decoder model for translating speech into the target text, and coarse-grained speech tokens, $C_{0}$; 2) Non-autoregressive acoustic model for acoustic details, $C_{0:16}$; 3) Codec model to convert discrete speech tokens back to the waveform. Abbreviation: S/A/I(Semantic/Acoustic/Isochrony Information), $C_{0}$/$C_{0:16}$(Codec layer 0/0-15), S/A-Enc(Semantic/Acoustic Encoder), ICM(Isochrony Control Module).
@@ 1 Introduction
To address these issues, we propose TransVIP, a speech-to-speech Translation framework with Voice and Isochrony Preservation. TransVIP employs a consecutive generation approach, simplifying the complex S2ST task into two sequential tasks while maintaining an end-to-end framework. The generation of the target speech is conditioned not only on the semantic information, as in conventional S2ST models, but also on the isochrony and acoustic information derived from the source speech. The corresponding overview of TransVIP is demonstrated in Figure 1. We evaluate the performance of TransVIP for French-English mutual translation using a subset of the CVSS-T test set [8]. The results demonstrate that TransVIP outperforms the publicly available SOTA models such as a larger Seamless Expressive model. The generated audio and video samples are available at https://aka.ms/transvip. The main contributions of the paper can be summarized as follows:
@@
    We introduce a framework for speech-to-speech translation tasks that employ a consecutive generation with joint inference. This method efficiently utilizes a variety of datasets through multi-task learning to overcome the challenge of scarce paired data during the training phase, while preserving the end-to-end nature during inference.
@@
    We propose to disentangle various information required to learn in the training stage through employing separated encoders. It can enhance the transfer of voice characteristics and isochrony temporal alignment from the source to the target speech in the translation process. Additionally, it facilitates the design of lightweight modules for more effective individual information learning.
@@ 2 Related Works
The task of speech quantization is to transform continuous speech features into discrete tokens. The quantization module is usually trained by self-supervised learning (SSL) methods. The speech tokens can be divided into two categories: semantic tokens and acoustic tokens[11]. The semantic tokens[12, 13, 14, 15, 16, 17] are usually learned by context prediction task. This kind of token is rich in context information and is suitable for downstream tasks like recognition. The acoustic tokens[18, 19, 20, 21, 22], also called the neural codec, is usually trained by reconstruction task through a discrete information bottleneck. This codec is suitable for audio compression and audio generation. SpeechTokenizer [9] combines the semantic token and acoustic token by semantic distillation. The latest works propose attribute disentanglement of neural codec for better generation task support. FAcodec[23] further disentangle codec into content, prosody, acoustic, and timbre through supervision and reversed gradient.
@@
With the recent advance in speech quantization and the strong in-context learning capability of the language model, the speech-language model has shown great potential in speech generation. The GSLM family[24, 25, 26] use semantic tokens to train language models on speech continuation. VoxtLM [27] integrates text tokens with semantic tokens to perform speech recognition, synthesis, and continuation in a single causal language model. AudioLM[11] uses both semantic tokens and acoustic tokens for high-quality speech continuation. VALL-E [10] extends this idea to zero-shot TTS and proposes auto-regressively generating the first layer and non-autoregressively the rest layers of the codec codes. Later work[28, 29, 30] have enhanced the functionality of zero-shot TTS systems, such as building cross-lingual synthesis and accurate alignment between text control signal and acoustic tokens.
@@
An end-to-end approach has been an ultimate goal for S2ST research. Some previous works[31, 32] have tried a direct end-to-end approach for S2ST. However, by far such methods still suffer great performance issues. A solution is to include semantic tokens(text, phoneme, or semantic speech tokens) as an intermediate result. Some recent works[33, 6, 34, 35] applied speech-to-unit translation(S2UT) that translate speech input into semantic speech tokens and then generate speech through a unit vocoder like Hifi-GAN. And UnitY[36] use separate speech-to-text and text-to-unit models that can be optimized together. However traditional vocoder based on semantic tokens like Hifi-GAN cannot generate high-quality speech. SeamlessExpressive[37] addresses this issue using a well-designed PRETSSEL vocoder. PolyVoice[38] and AudioPaLM[39] employees use two causal models to first translate speech into semantic tokens, then generate acoustic tokens based on semantic tokens, and then convert acoustic tokens to output speech. VioLA[40] and MSLM-S2ST[41] adopt a similar approach but use a single causal model in a multi-tasking way. However, these approaches perform several separate auto-regressive generations during the inference process. This, to a certain degree, contradicts the principle of end-to-end processing.
Figure 2: The illustration of the training framework of the Joint Enc-Dec Model. During the training, the losses from the target speech clip, i.e., a sub-part of the whole target speech, which serves as a prompt, are not aggregated when computing the Cross-Entropy (CE) loss. The corresponding codec labels are masked in the implementation. The semantic encoder and the auto-regressive decoder are initialized by a SeamlessM4T X2T model  [^2] https://ai.meta.com/blog/seamless-m4t. The semantic encoder is frozen during training. In inference, all the target speech input are replaced by source speech input.
@@ 3 TransVIP
3.1 Overview
Inspired by the recent advance in codec-based zero-shot TTS such as VALL-E [10], TransVIP contains three major parts: 1) A codec model for speech quantization and reconstruction, which is essential for auto-regressive speech generation (section 4.1). 2) An auto-regressive joint translation model to translate the input speech into coarse-grained speech (Section 3.1). 3) A textless non-autoregressive acoustic model to replenish acoustic details into the output speech (Section 4.2). The overall framework for generating target speech is shown in Figure 1.
In this section, we will present a detailed introduction to our joint translation model, which is the key component of our translation system. The training procedure of the joint model is illustrated in Figure 2. In section 4, we will discuss the codec and NAR model.
3.2 Consecutive Generation with Joint Inference
Directly generating target translation speech $Y$ from input speech $X$ presents a considerable challenge. Therefore, the primary objective of this model is to produce a coarse-grained speech $Y^{\prime}$ represented by a first-layer codec sequence. Instead of optimizing the direct probability distribution of $P(Y^{\prime}|X)$, our optimizing target is the joint probability $P(Y^{\prime},T|X)$ to preserve semantic accuracy in the text outputs alone, as $P(T|X)$. Hence, our approach focuses on modeling and maximizing the conditional expression:
@@ 3.2 Consecutive Generation with Joint Inference
3.3 Feature Disentanglement
In the real world, an ideal training dataset for speech-to-speech generation is not readily available, where the source and target speeches are roughly equal in length and uttered by the same speaker. A real dataset comprises paired speech, either spoken by different speakers or synthesized. Thus, we cannot guarantee voice and duration preservation with these corpora.
To address this issue, we decouple the input feature into three parts: semantic information($S$), acoustic information($A$), and isochrony information($I$). Formally, $X=(S,A,I)$. During training the inputs to the model are $S$ from source speech, and $(A,I)$ information from reference target speech. While during inference, $(A,I)$ are derived from source speech. In this way the input feature can always match the output speech in terms of voice characterizes $A$ and isochrony $I$.
@@ 3.3 Feature Disentanglement
$S$ is extracted using a pre-trained speech encoder from an encoder-decoder speech/text-to-text translation model, designed to provide minimal acoustic or speaker information. We also include the text encoder of the pre-trained model to enable text input and perform a knowledge distillation loss on the text output part. We freeze two pre-trained encoders during the training process.
@@
The acoustic information that we extract relates to the distinctive acoustic characteristics of a speaker’s voice. We assume that the translated text $T$ is independent of $A$. The relationship can be modeled as in Equation 3. We also take advantage of the one-direction mask in the auto-regressive decoder to model such a relation explicitly.
@@
We use an acoustic encoder, comprising transformer blocks that learn from scratch, to extract the acoustic information. This information is sum pooled along the time dimension into a single embedding, which then replaces the embedding of a separation token at the input of the auto-regressive decoder. This acoustic embedding will not affect text generation due to the causal property. It is compatible with both gradients passing during training and beam search during inference. In training, the input to the acoustic encoder is a speech prompt, i.e., a sub-part of the whole target speech. We zero out the loss in the same sub-part of the label to prevent information leakage. This combined with the sum pooling forms an information bottleneck to prevent complex semantic information from passing, allowing the encoder to learn meaningful acoustic information for voice preservation. Like Classifier-free Guidance, this replacement operation is not performed with a certain probability(in our experiment $p=0.5$) during training.
@@
We define a frame length, in our case 160ms, to quantify the duration of speech. Our Isochrony control module contains three embeddings: a regular position embedding, a reversed position embedding, and a voice activity detection (VAD) 0-1 embedding. The reversed position embedding provides the model with duration information on the number of additional tokens to be generated in each step of the inference process [42], and the VAD embedding identifies whether each frame is voice-active or not. Isochrony information is then represented by the sum of the three embedding sequences, which is then concatenated with the semantic feature $S$ before being fed into the decoder through cross-attention. This setup enables our model to align with both the total length and the voice-active regions of the input speech, which is especially beneficial for translating long speeches with pauses.
3.4 Multi-task Training
@@
  • S2ST Dataset comprises quadruple of ($X$, $T_{s}$, $T_{t}$, $Y$), where $T_{s}$ is the source text and $T_{t}$ is the target text. We either use $X$ or $T_{s}$ as input and consecutive $T_{t}$ and $Y$ as target labels, or we can reverse the roles of source input and target labels.
  • ST Dataset is a triplet of ($X$, $T_{s}$, $T_{t}$). In our model, the missing codec label does not impede the training of speech-to-text translation. We can use $X$ as the input and treat $T_{t}$ as an incomplete label that terminates with a separation token. In this setup, the isochrony feature is omitted and does not concatenate with the semantic feature. And, alternatively, $T_{t}$ can serve as the input with consecutive $X$ and $T_{s}$ as the label, mirroring the approach in the S2ST dataset.
  • ASR Dataset, which is highly accessible, consists of pairs ($X$, $T_{s}$). Unlike most other speech translation works, we use ASR data mainly for TTS training, fitting $P(Y^{\prime}|S,A,I,T)$. Here, $T_{s}$ is used as the input with consecutive $T_{s}$ and $X$ as the output labels, and the loss is not applied to the part of the label corresponding to $T_{s}$. We didn’t train on the recognition task because our semantic encoder is frozen, making this task ineffective.
4 Codec and Acoustic Modeling
@@ 4.1 Nerual Codec with Semantic Distillation
Our neural codec is designed to provide high reconstruction quality and compatibility with language model predictions. We refer to our codec as the multi-lingual Semantic-Aware Speech Codec (SASC), which integrates two existing methodologies: DAC [22] and SpeechTokenizer. DAC employs innovative factorized and L2-normalized codes in vector quantization and finely tuned hyperparameters to achieve state-of-the-art reconstruction quality. SpeechTokenizer, on the other hand, uses a Hubert encoder as a teacher and performs semantic distillation between it and the first-layer Residual Vector Quantization (RVQ) codes. This process enriches the RVQ codes with additional semantic information, facilitating downstream language modeling tasks.
@@
To enhance reconstruction quality, we have integrated the codec with factorized and L2-normalized codes. The discriminator, loss configuration, and training hyperparameters align with those used in DAC.
@@ 4.2 Textless Non-autoregressive Acoustic Modeling
Similar to previous work, the function of our non-autoregressive model is to enhance the acoustic details of generated speech by predicting the subsequent layers of Residual Vector Quantization (RVQ) codes based on the initial layer. We utilize a standard transformer equipped with adaptive layer normalization for this task. The model inputs the summation of embeddings from the first $n$ layers of RVQ codes to predict the $(n+1)$th layer, where $n=1,2,...,N$ and $N$ represent the total number of RVQ layers. A prompt is concatenated with the input embeddings along the time dimension. The prompt is the summation of full $N$ layers of embeddings of the prompt speech. During the training phase, the prompt and the part to be generated are in the same language. However, in the inference phase, they are in two distinct languages for the translation task. This creates a certain gap. Fortunately, our model trained using a multilingual dataset can effectively generalize and bridge this gap.
Our approach diverges from prior studies by excluding textual or phonetic inputs, rendering it a completely textless model. Previous research indicated that omitting semantic information such as text or phonemes typically significantly decreases intelligibility. However, our experiments demonstrate that the textless model either matches or surpasses the performance of models with text inputs, owing to semantic distillation within the SASC. A comparison of text and textless model is shown in appendix 6.3
The textless model offers two significant advantages: Firstly, it improves data accessibility, as the acoustic model can now be trained entirely with unsupervised data. Unlike supervised data, which requires aligned text labels and precise speech clip boundaries, unsupervised data can be arbitrarily segmented, effectively augmenting the dataset. Secondly, traditional non-autoregressive models that utilize phoneme inputs depend on accurate transcriptions to establish alignment and initiate generation. This requirement can be challenging, especially in zero-shot speech translation tasks where users might not provide transcriptions. Our textless model addresses this problem by eliminating the need for transcription.
@@
In previous work, the decoding of non-autoregressive models is typically performed greedily, as the decoding techniques used in auto-regressive models cannot be directly applied. Greedy decoding often leads to an "early decision error" problem where the model cannot revise previous outputs once generated. To address this, we propose a method called Layer Beam Search (LBS), which adapts the beam search algorithm for non-autoregressive models.
Consider a beam size of $B$ and a vocabulary size of $V$. In a typical beam search algorithm, a beam maintains $B$ hypotheses. At each decoding step, each hypothesis calculates the probability distribution over $V$, yielding $B\times V$ candidates. These candidates are then ranked by their aggregated scores, and the top $B$ are selected to form the new beam. This approach is impractical in non-autoregressive models due to the enormous vocabulary size at each decoding step. For example, for a speech segment containing $L$ tokens, and a codec codebook size of $C$, the effective vocabulary size becomes $|V|=C^{L}$, which is too large to enumerate.
To tackle this, we employ a sampling method to choose a reasonable number of candidates. Rather than sampling from the entire vocabulary space $V$, we sample from the top-K candidates at each token, thus creating a reduced vocabulary subspace $V^{\prime}$ of size $k^{L}$. We sample $N$ times from $V^{\prime}$ for each candidate in the original beam to form a total of $B*N$ new candidates. We then rank and form a new beam of size $B$ from these sampled candidates. The pseudo code for the Layer Beam Search is provided in appendix D.